OpenStudy (anonymous):

e^ { 2 x } - 6 e^ { x } +8 = 0 please help find the 2 solutions

OpenStudy (anonymous):

solve the quadratic equation \[z^2-6z+8=0\] for z and then replace z by \[e^x\]

OpenStudy (anonymous):

how did you get that quadratic equation?

OpenStudy (anonymous):

get \[(z-2)(z-4)=0\] \[z=2,z=4\] so \[e^x=4,x=\ln(4)\] or \[e^x=2,x=\ln(2)\]

OpenStudy (anonymous):

i get it because \[e^{2x}=(e^x)^2\]

OpenStudy (anonymous):

so i am thinking that this is \[(e^x)^2-6e^x+8\] which is a quadratic equation in \[e^x\]

OpenStudy (anonymous):

i think i understand how you got the equation now, thank you!